The selling price was $108. If the markup was 20 percent of the purchase price, what was the purchase price? What was the markup?
Answer by mukhopadhyay(490)

Assume purchase price as x;
x+20%x=108
=>x+.2x=108
=>1.2x=108
=>x=108/1.2=90
The markup was $108-$90=$18